OpenStudy (anonymous):

An electron is accelerated in a vacuum from rest through a potential difference of 850V. Show that the final momentum of the electron is 1.6 × 10–23 N s.

OpenStudy (anonymous):

The motion of an electron with that energy is nonrelativistic, so you can feel free to use \[ p = \sqrt{2mE } \] where \[ E = e\cdot \Delta V \] \[ e = 1.6\cdot 10^{-19} \text{ C} \] \[\Delta V = 850 \text{ V }\] \[ m = 9.1 \cdot 10^{-31} \text{ kg} \]